OpenStudy (compassionate):

Train A has a speed 15 mi/hr greater than train B. If train A travels 150 miles in the same time train B travels 120 miles, what are the speeds of the two trains?

OpenStudy (compassionate):

My logic being, \[x - 15 = B's \] Therefore \[\frac{ 150 }{ x } = \frac{ 120 }{ x - 15 }\]